UT 2019 09 Apr - 1st half SOUL / 2nd half TSIP-A.Conrad PI

OSA: D.G.Huerta

ISp: J. Morris

Mtn Manager: M. Wagner

ISA: B. Rothberg (Tucson)

AO: G. Taylor (Tucson)

Featured Players: J. Hill, J. Christou, and the ARCETRI AO team

All Problems - see G. Rahmer

UT 02:34 - getting ready to open

UT 02:42 - open - attempting to point and collimate

UT 03:24 - gopt seems to be working

-moving to auto-gain test

-will move to test sciene target at UT 04:00

UT 03:56 - testing offsets before trying a science target

UT 04:01 - moving to the AO Ref Star for the science field to see how the system responds, then will go to the science field which is 10" away.

UT 04:21 - turns out AO Ref star for test science field is a double star

UT 04:25 - running auto-gain on AO Ref Star

UT 04:34 - sending preset again - first try had interrupted system call error

-second time failed - multi-BCU error

UT 04:36 - sending preset again...

-techincal viewer shared buffer error (CCD47)

UT 04:38 - sending preset again

UT 04:43- still acquiring AO

-Rotator is at 301 (OK for NCPA)

UT 04:50 - taking first cube luci1.20190409.0069 - subframe was active - ignore

luci1.20190409.0070 - 2.51 sec exposure to test N30 fieldstop alignement - too faint

luci1.20190409.0071 - 10sec DIT - aligning N30 Fieldstop

luci1.20190409.0072 - confirmatory image

luci1.20190409.0073 - 2nd confirmatory image

luci1.20190409.0074 - first image

UT 05:04 - collecting data - preset presented an error not seen before - took time to figure out how to deal with it

luci1.20190409.0082 - AO loop opened offset was -45" -5" DET X Y - do not use this exposure

UT 05:14 - luci1.20190409.0083 - retaking -45" -5" DET X Y offset

UT 05:17 - return to main field OK - luci1.20190409.0084

UT 05:29 - paused - luci1.20190409.0092 - saw thicker clouds run overhead - AORef Star flux dropped - stopped just before the offset to PSF

UT 05:31 - resuming script - luci1.20190409.0093 - -45" -5" offset

UT 05:35 - luci1.20190409.0095 - cloud passed over - exposure may not be useful

UT 05:38 - AO loop opened - AO Ref star registering as R=17 (due to clouds)

-20 good exposures taken to date

UT 05:53 - clouds have dispersed - resuming - luci1.20190409.0096

UT 06:10 - Airmass is 1.20

UT 06:18 - rotator is 265.8

UT 06:39 - luci1.20190409.0123 - a few skip frames - exposure could be compromised

UT 06:57 - last exposure luci1.20190409.0133 - finished at airmass 1.336 rotator = 257

UT 06:58 - moving to a fainter star for testing

-- BarryRothberg - 09 Apr 2019
Topic revision: r1 - 09 Apr 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback